2.An A4 sheet of paper has length 29.69 ± 0.01 cm, width 21.1 ± 0.1 cm. Determine the area of the sheet of paper in cm2 with appropriate uncertainty.


1
Expert's answer
2021-05-30T13:28:32-0400
"A=(29.69)(21.1)=626.459\ cm^2\\
\frac{0.01}{29.69}+\frac{0.1}{21.1}=\frac{\Delta A}{626.459}\\
\Delta A=3\ cm^2\\A=(626\pm3)\ cm^2"